# Probabilities of the third type: Statistical Relational Learning and Reasoning with Relative Frequencies

### v3

@@ Abstract
Dependencies on the relative frequency of a state in the domain are common when modelling probabilistic dependencies on relational data. For instance, the likelihood of a school closure during an epidemic might depend on the proportion of infected pupils exceeding a threshold. Often, rather than depending on discrete thresholds, dependencies are continuous: for instance, the likelihood of any one mosquito bite transmitting an illness depends on the proportion of carrier mosquitoes. Current approaches usually only consider probabilities over possible worlds rather than over domain elements themselves. An exception are the recently introduced Lifted Bayesian Networks for Conditional Probability Logic, which express discrete dependencies on probabilistic data. We introduce functional lifted Bayesian networks, a formalism that explicitly incorporates continuous dependencies on relative frequencies into statistical relational artificial intelligence. and compare and contrast them with ifted Bayesian Networks for Conditional Probability Logic. Incorporating relative frequencies is not only beneficial to modelling; it also provides a more rigorous approach to learning problems where training and test or application domains have different sizes. To this end, we provide a representation of the asymptotic probability distributions induced by functional lifted Bayesian networks on domains of increasing sizes. Since that representation has well-understood scaling behaviour across domain sizes, it can be used to estimate parameters for a large domain consistently from randomly sampled subpopulations. Furthermore, we show that in parametric families of FLBN, convergence is uniform in the parameters, which ensures a meaningful dependence of the asymptotic probabilities on the parameters of the model.
@@ 1 Introduction
This distinction has first been methodically investigated in a strict logical famework by ? (?), who distinguished two concepts of probability, the “degree of confirmation”, which he calls “probability1”, and the “relative frequency (in the long run)”, which he calls “probability2”. Carnap goes on to formalise probability1 using a probability measure defined over so-called state descriptions, which we can identify as possible worlds in more modern terminology. Probability2 on the other hand is interpreted by the uniform measure on a given domain set itself.
Forty years later, in his seminal paper on the analysis of first-order logics of probability, Joseph Halpern (?) divided approaches to formalising probability in a relational setting along very similar lines. Halpern refers to logics encoding relative frequencies as Type I logics, while referring to logics encoding degrees of belief as Type II logics [^1] Somewhat unfortunately, this terminology is reverse to that of Carnap mentioned above. As a distinct category, Halpern also considers logics that combine both by expressing a degree of belief in a statement that mentions relative frequencies. He refers to those as Type III logics. Type III logics can express compound statements such as “With a likelihood of at least 10%, more than 60% of the population will have been ill by the end of the year.”
@@ 1.1 Queries Relating to Degrees of Belief vs Relative Frequency
Such trigger conditions that refer to relative frequencies are very common. (see (?, Table II) for further examples). This naturally also leads to Type III queries, in which the likelihood of a certain frequency event is adressed: “How likely is it that 1% of the population will be infected within the next four weeks?”
@@ 1.3 Related Work
The distribution semantics for PLP, first introduced in essence by ? (?) and then named and explicitly proposed as a semantics for probabilistic logic programming by ? (?) is given by a stratified logic program over independently distributed probabilistic ground facts on a domain $D$. More precisely, let $\rho\subseteq\sigma$ be signatures and let $\alpha_{i}\in[0,1]$ for atoms $R_{i}(\vec{x})$ from $\rho$. Let $\Pi$ be a stratified Datalog program with extensional signature $\rho$ and intensional signature $\sigma$. Then every $\rho$-structure $\mathfrak{Y}$ on $D$ induces a $\sigma$-structure $\mathfrak{X}$ on $D$ obtained by evaluating $\Pi$ with input $\mathfrak{Y}$. In this way, $\Pi$ lifts the probability distribution on $\rho$-structures given by making independent choices $R_{i}(\vec{x})$ of atoms from $\rho$ with probabilities $\alpha_{i}$,
@@
[…]
@@
The structure of the relational logistic regression network is subject to similar choices as the structure of the probabilistic logic program for the same scenario; we again choose to model a dependency of being considered social on having friedns, rather than vice versa. This implies the underlying DAG structure to be $\mathrm{Friends}(x,y)\longrightarrow\mathrm{Social}(x)$, with a single atomic formula $\mathrm{Friends(x,y)}$ annotating the node $\mathrm{Social}(x)$ and an associated positive weight $w$. The probability of an individual being social is then given by $\mathrm{sigmoid}(nw)$, where $n$ is the number of friends that individual has.
@@
? (?) introduced conditional probability logic (CPL) as a language for expressing Type I probabilities, and defined lifted Bayesian networks (LBN-CPL) in which the conditional probability tables depend on such CPL formulas. Therefore, LBN-CPL can clearly be indentified as a Type III formalism.
@@ 2.1 Conditional Probability Logic
We introduce some intutitive shorthands:
@@ 2.2 Lifted Bayesian Networks for Conditional Probability Logic
    For each $R\in\sigma$ a finite tuple of formulas ${\big{(}\chi_{R,i}(\vec{x})\big{)}}_{i\leq\nu_{R}}\in\mathrm{CPL}(\mathrm{par}(R))$, where $\mathrm{par}(R)$ is the signature of the $G$-parents of $R$, $\vec{x}$ is a sort-appropriate tuple of the correct length for $R$, such that $\forall\vec{x}\big{(}\bigvee_{i=1}^{\nu_{R}}\chi_{R,i}(\vec{x})\big{)}$ is valid (i.e. true in all $\mathrm{par}(R)$-structures) and if $i\neq j$ then $\exists\vec{x}\big{(}\chi_{R,i}(\vec{x})\wedge\chi_{R,j}(\vec{x})\big{)}$ is unsatisfiable. Such a set of $\chi_{R,i}$ is called a partition.
@@
    For each $R\in\sigma$ and each associated formula, a number denoted $\mu(R\ \mid\ \chi_{R,i})$ (or $\mu(R(\vec{x})\ \mid\ \chi_{R,i}(\vec{x}))$) in the interval $[0,1]$.
The semantics of lifted Bayesian networks are defined by grounding to a Bayesian network with respect to a given domain. So let $D$ be a finite domain. Then we can view a $\sigma$-structure with domain $D$ as a choice of truth value for each $R(\vec{a})$, where $R$ is a relation symbol in $\sigma$ and $\vec{a}$ is a tuple of elements of $D$ of the right length and the right sorts for $R$. Therefore, defining a probability distribution over the space of possible $\sigma$-structures with domain $D$ is equivalent to defining a joint probability distribution over the $R(\vec{a})$, viewed as binary random variables.
@@ 4.1 Expressivity of lifted Bayesian networks for conditional probability logic
This is an interesting parallel with PLP and shows that while LBN-CPL are defined using a graphical moel as a backbone, they could alternatively have been defined in the style of “independent probabilistic facts and determinsitic rules.” Note, however, that unlike in the case of PLP, the new relation $P(R_{i})$ that we introduce has just the same arity and sorts as $R$, and is not in the scope of any quantifier.
@@ 4.2 Expressivity of Functional Lifted Bayesian Networks
In this model, Whether a place actually meets the standard or not is stochastic, in that a certain proportion of establishments meet the required standarrds and some don’t. Similarly, there is a certain chance for any individual reviewer to be experienced. As these nodes do not have any parents, there are no further choices to be made beyond specifying their probabilities of occurrence.
@@
Finally, one could instead consider that regardless of how large the domain of possible inspectors, having only positive reviews will always be significantly different than at least one ispector noting concerns. To model this, we simply add the additional formula $\forall_{x}\mathrm{Positive\_review(x,y)}$. This will always evaluate to 0 or 1, and can be added as an additional factor with an additional weight $w_{2}$, so that $f\left(\alpha,\beta\right)=\mathrm{sigmoid}\left(w_{1}(\alpha-p_{o})+w_{2}\right)$.
@@ 5 Asymptotic Analysis of the Extrapolation Behaviour
A probability distribution $\mathbb{P}_{\infty}$ on $\Omega_{\infty}$ is the asymptotic limit of a lifted Bayesian network $\mathfrak{G}$ if for any sequence $D_{k}$ of domains which which tend uniformly to infinity and any generating set $A$ the limit of $\mathbb{P}_{\mathfrak{G},\vec{n}}(A)$ equals $\mathbb{P}_{\infty}(A)$. Two functional lifted Bayesian networks $\mathfrak{G}$ and $\mathfrak{G}^{\prime}$ are asymptotically equivalent if they share the same asymptotic limit $\mathbb{P}_{\infty}$.
@@
The condition that $f_{R}^{-1}\{0,1\}\subseteq{\{0,1\}}^{n}$ signifies that as long as among all of the parent formulas, there are both examples and counterexamples, the relation at the child node is not deterministic. This is important for asymptotic analysis since the behaviour of relations that occur with a very small but positive probability is radically different from that of relations whose probability is precisely 0. For the former, as long as the domain size is big enough, there will almost always be at least a few examples of this relation present among domain individuals; in the latter case, the relation will not occur no matter how large the domain.
@@
While ? (?) gives a stand-alone proof, we can derive it in a straightforward way from Theorem 5 and Proposition 1:
@@ 5.1 Proof of Theorem 4
For the proof we will use the classical model-theoretic notion of an extension axiom (?). the extensions axioms formalise the notion that any possible configuration of relations within a stucture do occur.
@@ 5.2 Examples
Consider a language with two sorts and two unary relation symbols $P$ and $R$, each with their own sort, and an underlying graph $P\longrightarrow R$. Let $P$ be annotated with the probability 0.5 and let $R$ be annotated with the aggregation function $f_{R}:[0,1]\rightarrow[0,1]$, where $f_{R}(a)=0.7$ if $a\geq 0.5$ and $f_{R}(a)=0.2$ otherwise. Note that this does not give rise to an FLBN by our definition because $f_{R}$ is not continuous. As the domain size increases, the probability that the relative frequency of $P$ among domain elements is at least 0.5 approaches 0.5. Therefore, asymptotically, the relative frequency of domain elements satisfying $R$ will either be close to 0.2 or close to 0.7, each with a probability of 0.5. This behaviour cannot be modelled by a quantifier-free lifted Bayesian network, as such a network could only specifiy a single probability number for the relation symbol $R$ (the only relation symbol mentioning its sort).
Another marked restriction of our framework is the omission of equality from our language. Indeed, including equality also causes the failure of asymptotic convergence. The counterexample relies on a famous result by (?), accessibly expounded in Spencer’s (?) monograph, that sparse random graphs with edge probabilities $(1/n)^{\alpha}$ with rational coefficients $\alpha$ do not have limit probabilities for first-order logic.
@@
Consider a two-sorted domain with a binary relation symbol $E$, with both entries of the first sort, and a unary relation symbol $R$ with one entry of the second sort. Consider the underlying graph $P\longrightarrow R$. Define an edge relation $E$ with the formula $x=y$ and the aggregation function $f_{E}:[0,1]\rightarrow[0,1]$, where $f_{E}(a)=\sqrt{(}a)$. This corresponds to a random graph in which edges are rawn with probability $(1/n)^{(1/2)}$. By the main result of (?), there is a first-order $\{E\}$-sentence $\varphi$ such that the probability that $\varphi$ holds in the random graph with $n$ nodes and edge probability $(1/n)^{(1/2)}$ diverges with increasing $n$. Now consider the aggregation function $f_{R}$ to be 1 if $\varphi$ holds and 0 otherwise. In this case, $R(x)$ will hold for either all or none of the elements of the domain, with the likelihood of both scenarios diverging as $n$ increases. This clearly rules out asymptotic convergence to a quantifier-free lifted Bayesian network, since in those the proportion of elements satisfying $R$ would converge to the probability $R$ is annotated with.
@@ 6 Further work
We now briefly consider the methods and algorithms that are available for learning LBN-CPL and FLBN. Let us consider LBN-CPL first. Here, parameter learning from interpretations amounts to learning the probabilities $\mu(R\ \mid\ \chi_{R,i})$ given one (or more) $\sigma$-structures. This case is closely aligned to parameter learning in other directed settings. In particular, in the case of a completely specified training model, the probabilities can be set to equal the frequencies encountered in the data. This can actually be expressed in terms of CPL itself: Set $\mu(R\ \mid\ \chi_{R,i})$ to be equal to the relative frequency $\left\|R(\vec{x})\mid\chi_{R,i}(\vec{x})\right\|_{\vec{x}}$. In the case of missing data, we can adapt the expectation-maximisation algorithm for learning parameters in ordinary Bayesian networks to CPL. Since we can view the LBN-CPL as a deterministic model over probabilistic facts (Proposition 1), this can be done by tying parameters known to be the same from the relational structure of the model and the formulas $\chi_{R,i}$ in just the same way as in probabilistic inductive logic programming (outlined for instance by (?, Section 8.5). This native support for parameter learning makes it particularly attractive for modelling tasks such as the ones outlined above. In the context of infectious disease dynamics for instance, the parameters have epidemiological meanings and their estimation from past data is an important part of informed decision making. While structure learning could in principle also be accomplished along the lines of the SLIPCOVER beam search algorithm (?), we anticipate that the significantly larger theory space resulting from the added probability quantifiers would make that difficult. Indeed, as far as we are aware even the problem of plain inductive logic programming in CPL without a Type III extension is yet to be addressed.

### v4

@@ Abstract
Dependencies on the relative frequency of a state in the domain are common when modelling probabilistic dependencies on relational data. For instance, the likelihood of a school closure during an epidemic might depend on the proportion of infected pupils exceeding a threshold. Often, rather than depending on discrete thresholds, dependencies are continuous: for instance, the likelihood of any one mosquito bite transmitting an illness depends on the proportion of carrier mosquitoes. Current approaches usually only consider probabilities over possible worlds rather than over domain elements themselves. An exception are the recently introduced Lifted Bayesian Networks for Conditional Probability Logic, which express discrete dependencies on probabilistic data. We introduce functional lifted Bayesian networks, a formalism that explicitly incorporates continuous dependencies on relative frequencies into statistical relational artificial intelligence, and compare and contrast them with lifted Bayesian networks for conditional probability logic. Incorporating relative frequencies is not only beneficial to modelling; it also provides a more rigorous approach to learning problems where training and test or application domains have different sizes. To this end, we provide a representation of the asymptotic probability distributions induced by functional lifted Bayesian networks on domains of increasing sizes. Since that representation has well-understood scaling behaviour across domain sizes, it can be used to estimate parameters for a large domain consistently from randomly sampled subpopulations. Furthermore, we show that in parametric families of FLBN, convergence is uniform in the parameters, which ensures a meaningful dependence of the asymptotic probabilities on the parameters of the model.
@@ 1 Introduction
This distinction has first been methodically investigated in a strict logical framework by ? (?), who distinguished two concepts of probability, the “degree of confirmation”, which he calls “probability1”, and the “relative frequency (in the long run)”, which he calls “probability2”. Carnap goes on to formalise probability1 using a probability measure defined over so-called state descriptions, which we can identify as possible worlds in more modern terminology. Probability2 on the other hand is interpreted by the uniform measure on a given domain set itself.
Forty years later, in his seminal paper on the analysis of first-order logics of probability, Joseph Halpern (?) divided approaches to formalising probability in a relational setting along very similar lines. Halpern refers to logics encoding relative frequencies as Type I logics, while referring to logics encoding degrees of belief as Type II logics [^1] Somewhat unfortunately, this terminology is reverse to that of Carnap mentioned above.. As a distinct category, Halpern also considers logics that combine both by expressing a degree of belief in a statement that mentions relative frequencies. He refers to those as Type III logics. Type III logics can express compound statements such as “With a likelihood of at least 10%, more than 60% of the population will have been ill by the end of the year.”
@@ 1.1 Queries Relating to Degrees of Belief vs Relative Frequency
Such trigger conditions that refer to relative frequencies are very common. (see (?, Table II) for further examples). This naturally also leads to Type III queries, in which the likelihood of a certain frequency event is addressed: “How likely is it that 1% of the population will be infected within the next four weeks?”
@@ 1.3 Related Work
The distribution semantics for PLP, first introduced in essence by ? (?) and then named and explicitly proposed as a semantics for probabilistic logic programming by ? (?) is given by a stratified logic program over independently distributed probabilistic ground facts on a domain $D$. More precisely, let $\rho\subseteq\sigma$ be signatures and let $\alpha_{i}\in[0,1]$ for atoms $R_{i}(\vec{x})$ from $\rho$. Let $\Pi$ be a stratified Datalog program with extensional signature $\rho$ and intensional signature $\sigma$. Then every $\rho$-structure $\mathfrak{Y}$ on $D$ induces a $\sigma$-structure $\mathfrak{X}$ on $D$ obtained by evaluating $\Pi$ with input $\mathfrak{Y}$. In this way, $\Pi$ lifts the probability distribution on $\rho$-structures given by making independent choices $R_{i}(\vec{x})$ of $\rho$-atoms with probabilities $\alpha_{i}$,
@@
along with probabilities $\alpha_{1}$ and $\alpha_{2}$ for the relation symbols $\mathrm{Friends}$ and $\mathrm{Makes\_social}$ respectively. Here, $\alpha_{1}$ sets the probability for any two individual to be social, and $\alpha_{2}$ models the probability of any individual friendship to cause a person to be considered social. These are considered to be independent causes for socialness, so if an individual has $n$ friends, then the probability for him to be social can be seen to be
@@
The structure of the relational logistic regression network is subject to similar choices as the structure of the probabilistic logic program for the same scenario; we again choose to model a dependency of being considered social on having friends, rather than vice versa. This implies the underlying DAG structure to be $\mathrm{Friends}(x,y)\longrightarrow\mathrm{Social}(x)$, with a single atomic formula $\mathrm{Friends(x,y)}$ annotating the node $\mathrm{Social}(x)$ and an associated positive weight $w$. The probability of an individual being social is then given by $\mathrm{sigmoid}(nw)$, where $n$ is the number of friends that individual has.
@@
? (?) introduced conditional probability logic (CPL) as a language for expressing Type I probabilities, and defined lifted Bayesian networks (LBN-CPL) in which the conditional probability tables depend on such CPL formulas. Therefore, LBN-CPL can clearly be identified as a Type III formalism.
@@ 2.1 Conditional Probability Logic
We introduce some intuitive shorthands:
@@ 2.2 Lifted Bayesian Networks for Conditional Probability Logic
    For each $R\in\sigma$ a finite tuple of formulas ${\big{(}\chi_{R,i}(\vec{x})\big{)}}_{i\leq\nu_{R}}\in\mathrm{CPL}(\mathrm{par}(R))$, where $\mathrm{par}(R)$ is the signature of the $G$-parents of $R$, $\vec{x}$ is a sort-appropriate tuple of the correct length for $R$, such that $\forall\vec{x}\big{(}\bigvee_{i=1}^{\nu_{R}}\chi_{R,i}(\vec{x})\big{)}$ is valid (i.e. true in all $\mathrm{par}(R)$-structures) and if $i\neq j$ then $\exists\vec{x}\big{(}\chi_{R,i}(\vec{x})\wedge\chi_{R,j}(\vec{x})\big{)}$ is unsatisfiable. Such a tuple is called a partition.
@@
    For each $R\in\sigma$ and each associated formula, a number denoted $\mu(R\ \mid\ \chi_{R,i})$ or $\mu(R(\vec{x})\ \mid\ \chi_{R,i}(\vec{x}))$ in the interval $[0,1]$.
The semantics of lifted Bayesian networks are defined by grounding to a Bayesian network with respect to a given domain. We can view a $\sigma$-structure with a finite domain $D$ as a choice of truth value for each $R(\vec{a})$, where $R$ is a relation symbol in $\sigma$ and $\vec{a}$ is a tuple of elements of $D$ of the right length and the right sorts for $R$. Therefore, defining a probability distribution over the space of possible $\sigma$-structures with domain $D$ is equivalent to defining a joint probability distribution over the $R(\vec{a})$, viewed as binary random variables.
@@ 4.1 Expressivity of Lifted Bayesian networks for Conditional Probability Logic
This is an interesting parallel with PLP and shows that while LBN-CPL are defined using a graphical model as a backbone, they could alternatively have been defined in the style of “independent probabilistic facts and determinsitic rules.” Note, however, that unlike in the case of PLP, the new relation $P(R_{i})$ that we introduce has just the same arity and sorts as $R$, and is not in the scope of any quantifier.
@@ 4.2 Expressivity of Functional Lifted Bayesian Networks
In this model, Whether a place actually meets the standard or not is stochastic, in that a certain proportion of establishments meet the required standards and some don’t. Similarly, there is a certain chance for any individual reviewer to be experienced. As these nodes do not have any parents, there are no further choices to be made beyond specifying their probabilities of occurrence.
@@
Finally, one could instead consider that regardless of how large the domain of possible inspectors, having only positive reviews will always be significantly different than at least one inspector noting concerns. To model this, we simply add the additional formula $\forall_{x}\mathrm{Positive\_review(x,y)}$. This will always evaluate to 0 or 1, and can be added as an additional factor with an additional weight $w_{2}$, so that $f\left(\alpha,\beta\right)=\mathrm{sigmoid}\left(w_{1}(\alpha-p_{o})+w_{2}\right)$.
@@ 5 Asymptotic Analysis of the Extrapolation Behaviour
A probability distribution $\mathbb{P}_{\infty}$ on $\Omega_{\infty}$ is the asymptotic limit of a lifted Bayesian network $\mathfrak{G}$ if for any sequence $D_{k}$ of domains which tend uniformly to infinity and any generating set $A$ the limit of $\mathbb{P}_{\mathfrak{G},\vec{n}}(A)$ equals $\mathbb{P}_{\infty}(A)$. Two functional lifted Bayesian networks $\mathfrak{G}$ and $\mathfrak{G}^{\prime}$ are asymptotically equivalent if they share the same asymptotic limit $\mathbb{P}_{\infty}$.
@@
The condition that $f_{R}^{-1}\{0,1\}\subseteq{\{0,1\}}^{n}$ signifies that as long as among all of the parent formulas, there are both examples and counterexamples, the relation at the child node is not deterministic. This is important for asymptotic analysis since the behaviour of relations that occur with a very small but positive probability is radically different from that of relations whose probability is precisely 0. For the former, as long as the domain size is big enough, there will almost always be at least a few examples of this relation present among domain individuals; in the latter case, the relation will not occur, no matter how large the domain.
@@
While ? (?) gives a stand-alone proof, we can derive it straightforwardly from Theorem 5 and Proposition 1:
@@ 5.1 Proof of Theorem 4
For the proof we will use the classical model-theoretic notion of an extension axiom (?). The extensions axioms formalise the notion that any possible configuration of relations within a structure do occur.
@@ 5.2 Examples
Consider a language with two sorts and two unary relation symbols $P$ and $R$, each with their own sort, and an underlying graph $P\longrightarrow R$. Let $P$ be annotated with the probability 0.5 and let $R$ be annotated with the aggregation function $f_{R}:[0,1]\rightarrow[0,1]$, where $f_{R}(a)=0.7$ if $a\geq 0.5$ and $f_{R}(a)=0.2$ otherwise. Note that this does not give rise to an FLBN by our definition because $f_{R}$ is not continuous. As the domain size increases, the probability that the relative frequency of $P$ among domain elements is at least 0.5 approaches 0.5. Therefore, asymptotically, the relative frequency of domain elements satisfying $R$ will either be close to 0.2 or close to 0.7, each with a probability of 0.5. This behaviour cannot be modelled by a quantifier-free lifted Bayesian network, as such a network could only specify a single probability number for the relation symbol $R$ (the only relation symbol mentioning its sort).
Another marked restriction of our framework is the omission of equality from our language. Indeed, including equality also causes the failure of asymptotic convergence. The counterexample relies on a famous result by ? (?), accessibly expounded in Spencer’s (?) monograph, that sparse random graphs with edge probabilities $(1/n)^{\alpha}$ with rational coefficients $\alpha$ do not have limit probabilities for first-order logic.
@@
Consider a two-sorted domain with a binary relation symbol $E$, with both entries of the first sort, and a unary relation symbol $R$ with one entry of the second sort. Consider the underlying graph $P\longrightarrow R$. Define an edge relation $E$ with the formula $x=y$ and the aggregation function $f_{E}:[0,1]\rightarrow[0,1]$, where $f_{E}(a)=\sqrt{(}a)$. This corresponds to a random graph in which edges are drawn with probability $(1/n)^{(1/2)}$. By the main result of ? (?), there is a first-order $\{E\}$-sentence $\varphi$ such that the probability that $\varphi$ holds in the random graph with $n$ nodes and edge probability $(1/n)^{(1/2)}$ diverges with increasing $n$. Now consider the aggregation function $f_{R}$ to be 1 if $\varphi$ holds and 0 otherwise. In this case, $R(x)$ will hold for either all or none of the elements of the domain, with the likelihood of both scenarios diverging as $n$ increases. This clearly rules out asymptotic convergence to a quantifier-free lifted Bayesian network, since in those the proportion of elements satisfying $R$ would converge to the probability $R$ is annotated with.
@@ 6 Further Work
We now briefly consider the methods and algorithms that are available for learning LBN-CPL and FLBN. Let us consider LBN-CPL first. Here, parameter learning from interpretations amounts to learning the probabilities $\mu(R\ \mid\ \chi_{R,i})$ given one (or more) $\sigma$-structures. This case is closely aligned to parameter learning in other directed settings. In particular, in the case of a completely specified training model, the probabilities can be set to equal the frequencies encountered in the data. This can actually be expressed in terms of CPL itself: Set $\mu(R\ \mid\ \chi_{R,i})$ to be equal to the relative frequency $\left\|R(\vec{x})\mid\chi_{R,i}(\vec{x})\right\|_{\vec{x}}$. In the case of missing data, we can adapt the expectation-maximisation algorithm for learning parameters in ordinary Bayesian networks to CPL. Since we can view the LBN-CPL as a deterministic model over probabilistic facts (Proposition 1), this can be done by tying parameters known to be the same from the relational structure of the model and the formulas $\chi_{R,i}$ in just the same way as in probabilistic inductive logic programming (outlined for instance by ? (?, Section 8.5)). This native support for parameter learning makes it particularly attractive for modelling tasks such as the ones outlined above. In the context of infectious disease dynamics for instance, the parameters have epidemiological meanings and their estimation from past data is an important part of informed decision-making. While structure learning could in principle also be accomplished along the lines of the SLIPCOVER beam search algorithm (?), we anticipate that the significantly larger theory space resulting from the added probability quantifiers would make that difficult. Indeed, as far as we are aware even the problem of plain inductive logic programming in CPL without a Type III extension is yet to be addressed.